A boat travels 1 km upstream and 1km back. The time for the round trip is 10 hours. The speed of the stream is 2 km/hr. What is the speed of the boat in still water?
Answer by scott8148(5891)

let x=speed of boat in still water ___ time=distance/rate

(1/(x+2))+(1/(x-2))=10 ___ (x-2)+(x+2)=10(x-2)(x+2) ___ 4x=10x^2-40

0=5x^2-2x-20